You may want to reference (pages 754 - 765) sectio
Answer #1

A) pH=pka-log Salt] [salt [0.105] -4.76-log (0.1651 -4.56 3) brlog base] log [0240] =3.36-10g0115 o.113 3.68 pH-14-pOH pH-10.